How would I express 5x-2=6y in standard form? I am trying to help my 13 year old 7th grader who is taking 9th grade algebra I! I LOVE math, but "Mom doesn't know what she is talking about!"
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi,
*Note: standard form for an equation of a line is: Ax +By = C
where A > 0 and, if possible, A, B, and C are relatively prime integers.
5x-2 = 6y Putting into standard form
subtracting 6y from both sides and adding 2 to both sides of the equation.
5x - 6y = 2

Answer by Alan3354(69443)

Standard form is Ax + By = C, all integers
5x-2 = 6y
Add2
5x = 6y + 2
Subtract 6y
5x - 6y = 2